OpenStudy (anonymous):

(2 tan 30 degree)^2 + (csc 315 degree)^2 - sin 150 degree= what?

OpenStudy (anonymous):

=9/4+[2]-1/2 =(9+8-4)/2 =13/2

OpenStudy (anonymous):

i dont have that as one of my answer choices

OpenStudy (anonymous):

let me check my math...doing this in my head

OpenStudy (anonymous):

ah...ic my error

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

2.83...I guess you need a calculator for this one

OpenStudy (anonymous):

good now?

OpenStudy (anonymous):

my answer choices are -1/2 -7/6 17/6 7/2 -1

OpenStudy (anonymous):

17/6

OpenStudy (anonymous):

same thing